OpenStudy (anonymous):

Determine if the function is even, odd or neither f(x)=-3x^2+4

OpenStudy (anonymous):

Even

OpenStudy (anonymous):

can you show your work please

OpenStudy (anonymous):

AND THIS ?? f(x)=2x^3−4x

OpenStudy (anonymous):

f(x) = f(-x) ? -3x^2+4 = -3(-x)^2+4 (-x)^2=x^2 because (-x)(-x)=x^2

OpenStudy (anonymous):

To check if a function is even, substitute in -x for x and see if the resulting function is the same as the original

OpenStudy (anonymous):

CAN YOU DO THE SECOND PLEASE??

OpenStudy (anonymous):

f(x)=2x^3−4x is odd

OpenStudy (anonymous):

CAN YOU SHOW ME THE STEPS @larryboxaplenty

OpenStudy (anonymous):

what happens to the -3 in the first question?

OpenStudy (anonymous):

-f(x) = f(-x) ? -(2x^3-4x) = -2x^3+4x = f(-x) = 2(-x)^3-4(-x) = -2x^3+4x

OpenStudy (anonymous):

ok thanks. but can you answer my question above ^

OpenStudy (anonymous):

Even function: f(x) = f(-x) ? Odd function -f(x) = f(-x)?

OpenStudy (anonymous):

f(x)=2x^3−3x^2−4x+4??

OpenStudy (anonymous):

@larryboxaplenty

OpenStudy (anonymous):

try that one please

OpenStudy (anonymous):

Nether

OpenStudy (anonymous):

steps please :)

OpenStudy (anonymous):

If you look at the equation, you can see it,

OpenStudy (anonymous):

@larryboxaplenty what do you mean?

OpenStudy (anonymous):

I dont have to solve this to know its neither??

OpenStudy (anonymous):

@larryboxaplenty

OpenStudy (anonymous):

you have to see that its not even nor odd